OpenStudy (anonymous):

What is the y-intercept for this linear equation? 6x-y=9

OpenStudy (anonymous):

-9 Why? Change opriginal equation into slope-intercept form: y=mx+b and b is the y-intercept... So solve for y 6x-y=9 (subtract 6x to both sides) -y=-6x+9 (get rid of the negative attached to the y by distributing or multiplying the -1 to all things)... y=6x-9... So its -9 cause its in y=mx+b form. Does that make sense?